OpenStudy (anonymous):

At room temperature, the surface tension of water is 72.0 mJ·m–2. What is the energy required to change a spherical drop of water with a diameter of 1.20 mm to five smaller spherical drops of equal size? The surface area of a sphere of radius r is 4π r2 and the volume is 4π r3/3

OpenStudy (anonymous):

first you need to know that surface tension is energy per unit area. So initial energy =Ei = 72 * 4pi R^2 R= 0.6 mm final energy = Ef = 72 * 5 * 4 pi r^2 where r is the radius of the smaller drops You may find r by equating initial and final volumes 4/3 pi R^3 = 5 * 4/3 pi *r^3